OpenStudy (anonymous):

I need a little bit of help, they are more like true/false questions, so i'll post them (:

OpenStudy (anonymous):

\[a^{2/3} = \frac{ a^{2} }{ a^{3} }\]

OpenStudy (lime):

a^(2/3) = (a2/a3) a^(2/3) = (1/a) a = 1 I'm assuming the answer is true/positive.

OpenStudy (anonymous):

how did you get 1?

OpenStudy (lime):

a^(5/3) is the improper fraction of a^(2/3). Divide 5/3. What do you get for a mixed fraction?

OpenStudy (anonymous):

1 2/3

OpenStudy (lime):

1 2/3. a^(2/3) = (a2/a3) Replace the '(a2/a3)' a's with 1's.

OpenStudy (anonymous):

ok.. i get it

OpenStudy (lime):

The solution is 1. Unless there is something left out of the equation, that's it. :)

OpenStudy (anonymous):

alright